find the sum of each.
2.3+2.6+2.9+....; 40th term

Answer by htmentor(1343)

This is an arithmetic sequence with common difference = 0.3
The nth term of this sequence can be written a(n) = a + 0.3*(n-1), where a = the 1st term
So a(40) = 2.3 + 0.3(39) = 14
The sum of the first 40 terms is S(40) = (40/2)(a + a(40)) = 20(2.3+14) = 326